OpenStudy (anonymous):

I just need some help on this problem I been stuck on. Find the distance between parallel lines w/the equations 3x-5y=1 and 3x-5y=-3

OpenStudy (anonymous):

The distance between a point (u,v) and a line a x + by + c = 0 is \[ d=\frac{ |a u + b v +c | }{\sqrt{ a^2 + b^2}} \] Take a point (u,v) on the first line and apply this formula. You can take (u,v)=(2,1) for example

OpenStudy (anonymous):

Write the lines as 3x-5y-1=0 and 3x-5y+3=0

OpenStudy (anonymous):

\[ d=\frac{ |3(2) +(-5)(1) +3 | }{\sqrt{ 3^2 +5^2}}= \frac{ 4 }{\sqrt{34}}= \]

OpenStudy (anonymous):

ohhh thank you I understand now

OpenStudy (anonymous):

one more problem i been stuck on find equations of the lines that bisect the acute and obtuse angles formed by the graphs of 6x + 8y = -5 and 2x - 3y = 4

OpenStudy (anonymous):

Post is as a different problem and let me know about ti

OpenStudy (anonymous):

will do